OpenStudy (anonymous):

help please x and y intercepts of y=-|x+10|

OpenStudy (anonymous):

for x intercept 0=-|x+10| right ?

OpenStudy (anonymous):

0=-x+10

OpenStudy (anonymous):

y int y=10 x int x=10, -10

OpenStudy (haseeb96):

first solve for y intercept u know that for y intercept we put x=0 in the equation so here y=-Ix+10I y=-x-10 now put x=0 y=-10 -10 is yur y intercept and again put y=0 for x intercept y=-Ix+10I y=-x-10 y=o put in the above equation 0=-x-10 x=-10 this is yur x intercept